# Ultra-wideband Time Difference of Arrival Indoor Localization: From Sensor Placement to System Evaluation

### v1

@@ Abstract
Wireless indoor localization has attracted significant research interest due to its high accuracy, low cost, lightweight design, and low power consumption. Specifically, ultra-wideband (UWB) time difference of arrival (TDOA)-based localization has emerged as a scalable positioning solution for mobile robots, consumer electronics, and wearable devices, featuring good accuracy and reliability. While UWB TDOA-based localization systems rely on the deployment of UWB radio sensors as positioning landmarks, existing works often assume these placements are predetermined or study the sensor placement problem alone without evaluating it in practical scenarios. In this article, we bridge this gap by approaching the UWB TDOA localization from a system-level perspective, integrating sensor placement as a key component and conducting practical evaluation in real-world scenarios. Through extensive real-world experiments, we demonstrate the accuracy and robustness of our localization system, comparing its performance to the theoretical lower bounds. Using a challenging multi-room environment as a case study, we illustrate the full system construction process, from sensor placement optimization to real-world deployment. Our evaluation, comprising a cumulative total of $39$ minutes of real-world experiments involving up to five agents and covering $2608$ meters across four distinct scenarios, provides valuable insights and guidelines for constructing UWB TDOA localization systems.
Index Terms:
Localization, Sensor Fusion, Hardware-Software Integration in Robotics
@@ I Introduction
Indoor localization technology is rapidly advancing, offering a promising future where accurate positioning enables a broad spectrum of applications, including robotics, virtual/augmented reality (VR/AR), and seamless navigation with location-based services. Precise localization and ubiquitous communication have become essential expectations for the sixth generation (6G) wireless systems [1]. A conceptual diagram demonstrating the indoor positioning service in a shopping mall is shown in Figure 1. For indoor robotics applications, visual-inertial odometry (VIO) and visual simultaneous localization and mapping (SLAM) are commonly employed techniques to achieve precise 6 degrees-of-freedom localization, with cameras providing the main source of information. However, visual localization algorithms might face challenges under dynamic lighting conditions and partial or temporary occlusion of the cameras. More importantly, capturing raw images of the surrounding environment will pose challenges for security and privacy [2].
@@
Compact and computationally-constrained indoor robots and smart devices have led researchers to pursue localization methods leveraging low-power and lightweight sensors. Ultra-wideband (UWB) radio technology has been shown to provide potential high-accuracy time of arrival (TOA) measurements with low power consumption compared to commonly used sensors such as camera, radar, and light detection and ranging (LiDAR) in the field of robotics. UWB chips have been integrated into the latest generations of consumer electronics including smartphones and smartwatches to support low-latency spatially-aware interactions [3].
@@
Figure 1: A conceptual diagram demonstrating the deployment of an indoor localization system in a shopping mall. The image illustrates heterogeneous agents, including ground robots and flying robots, leveraging the indoor positioning system to navigate seamlessly alongside customers while providing various services, such as item delivery or cleaning service. (This conceptual image is created from an image we generated using DALL·E 3 text-to-image models developed by Open AI.)
@@
Similar to the Global Positioning System (GPS) [4], a UWB-based positioning system requires UWB radios (also called anchors) to be pre-installed in the environment as a constellation with known positions, which in turn serve as landmarks for positioning. In robotics, the two main ranging schemes used for UWB localization are (i) two-way ranging (TWR) and (ii) time difference of arrival (TDOA) (see Figure 2). In TWR, the UWB module mounted on the robot (also called the tag) communicates with an anchor and acquires range measurements through two-way communication. In TDOA, UWB tags compute the difference between the arrival times of the radio packets from two anchors as TDOA measurements. Compared with TWR, TDOA does not require active two-way communication between an anchor and a tag, thus enabling localization of a large number of devices. This scalability is crucial for scenarios that require localizing numerous agents simultaneously, such as providing positioning service for customers in large shopping malls and coordinating hundreds of drones for indoor drone shows. Therefore, we focus on developing UWB TDOA-based localization systems to facilitate the coordination of a large number of agents.
Considerable attention has been devoted to mitigating non-line-of-sight (NLOS) and multi-path radio propagation to enhance UWB-aided localization performance. However, these studies often assume the placement of UWB anchors is predetermined and overlook its critical impact on the localization performance. Existing sensor placement studies [5] primarily focus on finding a theoretically optimal sensor placement solution under restrictive assumptions, which often encounter challenges when applied to real-world cluttered scenarios. In our previous work [6], we proposed a more realistic optimal sensor placement algorithm that considers obstacles and multi-room scenarios. However, we only conducted simulation and static experiments in a lab setup without evaluating the algorithm in practical scenarios. In this article, we fill this gap by approaching the UWB TDOA localization from a system-level perspective, integrating sensor placement as a key component and conducting practical evaluation in real-world environments. Using pedestrian tracking problem as an example, we fuse UWB TDOA measurements with a low-cost inertial measurement unit (IMU) sensor to achieve 6 degrees-of-freedom pose estimates. We employ an error-state Kalman filter (ESKF), favored for its widespread use and effectiveness in practical applications, and compare our results against the theoretical lower bound. Using a challenging multi-room environment as a case study, we demonstrate the entire process of setting up a UWB TDOA localization system, from optimizing UWB anchor placement to deploying and testing the system in a real-world environment. To the best of our knowledge, this comprehensive evaluation, presenting a UWB TDOA-based localization system from a system-level perspective, has not been shown in the literature. A video summarizing our experiment process is available at http://tiny.cc/uwb_tdoa_ram24. Our main contributions can be summarized as follows:
  • We demonstrate the accuracy and robustness of the localization system through extensive real-world experiments and compare its performance to theoretical lower bounds.
  • We present the entire process of setting up a UWB TDOA localization system in a real-world multi-room environment, from sensor placement optimization to system deployment and evaluation against the theoretical analysis.
  • We will release the datasets collected during our experiments upon the acceptance of this work to benefit the research community.
II Related Work
As UWB measurements, like any other radio frequency (RF) signals, are often affected by obstacle-induced NLOS and multi-path radio propagation, multiple approaches have been proposed to improve localization accuracy under measurement corruption. M-estimators [7] are commonly used as versatile tools to mitigate the influence of measurement outliers by employing robust cost functions to downweight the impact of large measurement residuals. Researchers have also explored novel models to represent UWB measurement residuals including both parametric [8] and non-parametric models, such as neural networks [9], Gaussian processes [10], and Gaussian mixture models [11], aiming to enhance localization performance. Furthermore, continuous-time estimation techniques [12, 13] have garnered attention and have been investigated for their applicability in asynchronous UWB-aided localization systems.
@@ II Related Work
The anchor-tag geometry is widely recognized as a significant contributing factor to the performance of TDOA-based localization. Consequently, ensuring an optimized design for the placement of UWB anchors is crucial to guarantee reliable and accurate positioning performance. The sensor placement problem has most notably been studied by the GPS community. Geometric dilution of precision (GDOP) is a classical performance metric widely used to evaluate the quality of a GPS satellite configuration [14]. Cramér-Rao lower bound (CRLB) [15] is a more general performance metric for sensor placement, which considers statistical properties of sensor measurements to evaluate estimation performance. In our previous work [6], we extended the optimal sensor placement analysis to balance the effects of anchor-tag geometry and NLOS measurement biases in cluttered indoor environments. However, our experimental validation process was limited to static experiments conducted in a laboratory setup, using multilateration as the position estimator.
In this article, we present a UWB TDOA localization system approached from a system-level perspective, integrating sensor placement as a crucial component and evaluating it through practical experiments in real-world scenarios. To achieve 6 degrees-of-freedom poses while maintaining a low-power and lightweight design, we fuse UWB TDOA measurements with a low-cost IMU sensor through an error-state Kalman filter (ESKF) [16]. The integration and evaluation of this system in real-world settings demonstrate the impact of sensor placement within a complete localization framework. Through a comprehensive case study in a challenging multi-room environment, we illustrate the entire process of system construction, from sensor placement optimization to real-world deployment. We evaluate the localization performance by comparing it to the corresponding theoretical analysis, which provides valuable insights and guidelines for developing UWB TDOA localization systems.
@@ III-B Sensor Placement Optimization
Sensor placement optimization is often conducted through the lens of statistical parameter estimation, which optimizes the performance bounds for estimating the parameters of interest with a given model. Considering the UWB TDOA localization system in particular, we analyze the lower bound on the position estimate based on the UWB anchor placement and the TDOA measurement model. This analysis provides insights into the fundamental limitations of the system’s localization accuracy. In our previous work [6], we leverage the mean-squared error (MSE) metric to evaluate the UWB TDOA localization performance for a region of interest in cluttered environments. The estimated tag position $\hat{\mathbf{p}}$ is characterized by its bias $\textrm{Bias}(\hat{\mathbf{p}})$ and covariance matrix $\textrm{Cov}(\hat{\mathbf{p}})$. The MSE of an estimate $\hat{\mathbf{p}}$ of the true value $\mathbf{p}$ can be decomposed as
@@
Optimal anchor placement is essential for ensuring reliable localization performance in UWB TDOA localization systems, particularly in cluttered and geometrically challenging environments. Therefore, it is important to optimize the anchor positions for regions or trajectories of interest during system design. We define a set $\mathcal{A}$ containing all possible anchor configurations with $\bm{\alpha}\in\mathcal{A}$ and use a block coordinate-wise minimization (BCM) algorithm to find the optimal placement of UWB anchors $\bm{\alpha}^{\star}$ that minimizes $\mathcal{M}(\bm{\alpha})$:
@@ III-C UWB TDOA-Inertial Navigation System
To facilitate the localization of multiple heterogeneous agents, the motion model used in this work is a 3D kinematic motion model with IMU measurements as inputs [18]. This is generally referred to as an inertial navigation system (INS). The relevant equations and a detailed description of the motion model can be found in [18].
Figure 4: The multi-agent pedestrian localization experiment with five agents in constellation $1$ is shown in the photo above. The estimated together with the ground truth trajectories of each agent during the experiments are shown in the bottom plots. Readers are encouraged to view our supplementary video (http://tiny.cc/uwb_tdoa_ram24) to gain a better insight into our experimental process and evaluate the robustness of the localization performance.
The UWB TDOA measurements complement the motion dynamics by providing drift-free difference-of-distance measurements between UWB anchor pairs and the UWB tag. The measurement model for an anchor pair $\{\bm{a}_{i},\bm{a}_{j}\}$ is described as follows:
@@
To estimate the system state, we select the ESKF [16] algorithm for its proven effectiveness and widespread use in practical applications. In this formulation, inertial dead reckoning is used to propagate the state (4) forward in time using a 3D kinematic motion model with the IMU measurements as inputs. The uncertainty associated with dead reckoning is estimated in the prediction step. In the correction step, the error in the predicted state and the corresponding uncertainty are estimated using the UWB measurements. The error state corresponding to the state (4) is
@@
where $\delta\mathbf{p}$ and $\delta\mathbf{v}$ represent the errors in IMU position and linear velocity, respectively. Errors in the accelerometer and gyroscope biases are represented by $\delta\mathbf{b}_{a}$ and $\delta\mathbf{b}_{\omega}$, respectively. The local angular error $\delta\bm{\theta}$ is related to small differential rotations $\delta\mathbf{q}$ by: $\delta\mathbf{q}=\left(1,~{}\frac{1}{2}\delta\bm{\theta}\right)$, $|\delta\bm{\theta}|\ll 1$.
The correction step estimates the error between the dead reckoned state and the state consistent with UWB TDOA measurements. The estimated error (6) is then composed with the state to compensate for the accumulated drift: $\hat{\mathbf{x}}=\mathbf{x}\oplus{\delta\mathbf{x}}$, where $\oplus$ is a generic composition operator which represents (i) quaternion multiplication for the orientation error, $\mathbf{q}=\mathbf{q}\otimes\delta\mathbf{q}$, and (ii) the addition operation for the remaining states, $\mathbf{p}=\mathbf{p}+\delta\mathbf{p}$. We refer the reader to [19] for a description of the prediction and correction step of the ESKF.
@@ III-D System Architecture and Deployment Details
To evaluate localization performance, we assembled five UWB handheld devices, as shown in Figure 3d, specifically designed for multi-agent localization purposes. We leveraged the Loco Positioning System (LPS) from Bitcraze, which employs low-cost DWM1000 UWB radios, to construct our UWB TDOA localization system. To maintain cost-effectiveness, each handheld device is equipped with an inexpensive STMicro LSM9DS1 IMU and a DWM1000 UWB tag. The onboard computer for each device is a Raspberry Pi 4 Model B with 8GB RAM, paired with a power bank serving as the power supply. These elements are assembled within a 3D-printed case for portability. The total cost for a baseline system, including eight LPS UWB anchors from Bitcraze and one handheld device, falls below $2000$ USD using commercially available components, which can be further reduced by using raw sensor modules.
To enhance the portability of the localization system, all hardware components are configured as plug-and-play, including both the anchors and the handheld devices. The onboard computer on each handheld device runs an Ubuntu 20.04 LTS operating system, using Robot Operating System (ROS) Noetic as the middleware. The ESKF localization algorithm is implemented in C++ and runs at a frequency of 420 Hz, consuming approximately $35\%$ of a single CPU core and $0.4\%$ of the onboard computer’s RAM.
During the system deployment, we use a millimeter-accurate Leica total station (see Figure 3a) to establish the inertial frame and survey each anchor’s position in this frame. Utilizing the “Orientate to line” method in the Leica Captivate surveying field software, we surveyed two points to define the “East axis”, with the first serving as the origin and the second along the x-axis of the inertial frame we aim to define. After establishing the inertial frame, we conducted a survey of the pre-installed UWB anchors, measuring their 3D positions within this inertial frame with millimeter-level accuracy. Our UWB localization system provides drift-free positioning relative to the established inertial frame.
@@ IV Experiments
To fully demonstrate the localization performance of our UWB TDOA localization system, we conducted extensive experiments in a variety of real-world environments. We first demonstrate the multi-agent pedestrian localization performance and compare the localization accuracy resulting from two different anchor placements. Then we showcase the system’s ability to provide reliable positioning services under sensor occlusion. Furthermore, we demonstrate how sensor placement analysis effectively evaluates the degradation in positioning performance caused by the challenging anchor-tag geometry in a staircase scenario. Finally, using a challenging multi-room environment composed of a cafeteria and a narrow hallway as a case study, we illustrate the entire construction process of our UWB TDOA localization system, from sensor placement optimization to real-world deployment, and evaluate the localization performance compared to the theoretical lower bounds. A video summarizing our experiment process is available at http://tiny.cc/uwb_tdoa_ram24.
Figure 5: Comparison of the average experimental root-mean-square error (RMSE) for multi-agent pedestrian localization in Const. $1$ and $2$, along with localization under occlusion in Const. $1$ (Const. 1 Backpack), shown as blue bars, against the theoretical RMSE lower bounds represented by green bars.

### v2

# Wenda Zhao    Abhishek Goudar    Mingliang Tang    Angela P. Schoellig

@@ Abstract
Ultra-wideband (UWB) time difference of arrival (TDOA)-based localization has emerged as a scalable positioning solution for mobile robots, consumer electronics, and wearable devices, featuring good accuracy and reliability. While UWB TDOA-based localization systems rely on the deployment of UWB radio sensors as positioning landmarks, existing works often assume these placements are predetermined or study the sensor placement problem alone without evaluating it in practical scenarios. In this article, we bridge this gap by approaching the UWB TDOA localization from a system-level perspective, integrating sensor placement as a key component and conducting practical evaluation in real-world scenarios. Through extensive real-world experiments, we demonstrate the accuracy and robustness of our localization system, comparing its performance to the theoretical lower bounds. Using a challenging multi-room environment as a case study, we illustrate the full system construction process, from sensor placement optimization to real-world deployment. Our evaluation, comprising a cumulative total of $39$ minutes of real-world experiments involving up to five agents and covering $2608$ meters across four distinct scenarios, provides valuable insights and guidelines for constructing UWB TDOA localization systems.
@@ I Introduction
Indoor localization technology is rapidly advancing, offering a promising future where accurate positioning enables a broad spectrum of applications, including robotics, virtual/augmented reality (VR/AR), and seamless navigation with location-based services. Precise localization and ubiquitous communication have become essential expectations for the sixth generation (6G) wireless systems [18, 9]. A conceptual diagram demonstrating the indoor positioning service in a shopping mall is shown in Figure 1. For indoor robotics applications, visual-inertial odometry (VIO) and visual simultaneous localization and mapping (SLAM) are commonly employed techniques to achieve precise 6 degrees-of-freedom localization, with cameras providing the main source of information. However, visual localization algorithms might face challenges under dynamic lighting conditions and partial or temporary occlusion of the cameras. More importantly, capturing raw images of the surrounding environment will pose challenges for security and privacy [16].
@@
Compact and computationally-constrained indoor robots and smart devices have led researchers to pursue localization methods leveraging low-power and lightweight sensors. Ultra-wideband (UWB) radio technology has been shown to provide potential high-accuracy time of arrival (TOA) measurements with low power consumption compared to commonly used sensors such as camera, radar, and LiDAR in the field of robotics. UWB chips have been integrated into the latest generations of consumer electronics to support low-latency spatially-aware interactions [1, 11].
@@
Figure 1: A conceptual diagram demonstrating the deployment of an indoor localization system in a shopping mall. The image illustrates heterogeneous agents, including ground robots and flying robots, leveraging the indoor positioning system to navigate seamlessly alongside customers while providing various services. (This conceptual image is created from an image we generated using DALL·E 3 text-to-image models developed by Open AI.)
Similar to the Global Positioning System (GPS) [4], UWB-based positioning systems rely on pre-installed anchors with known locations that serve as environmental landmarks. In robotics, UWB localization typically employs either two-way ranging (TWR) or time difference of arrival (TDOA) (Figure 2). TWR estimates distances through two-way communication between a tag and an anchor, whereas TDOA computes the arrival-time differences of signals from anchor pairs. Unlike TWR, TDOA does not require active two-way communication, enabling scalable localization of a large number of devices. This scalability is critical for applications involving many agents, such as large-scale indoor navigation and coordinated multi-robot systems. Therefore, we focus on UWB TDOA-based localization in this work.
Significant effort has been devoted to mitigating non-line-of-sight (NLOS) and multipath effects to improve UWB-based localization. However, most existing approaches assume fixed anchor placement and overlook its critical influence on localization performance. Prior sensor placement studies [8] often rely on restrictive assumptions that limit their applicability in cluttered real-world environments. In our previous work [22], we proposed a more realistic sensor placement method that accounts for obstacles and multi-room scenarios, but validation was limited to simulation and controlled laboratory settings.
This article addresses this gap by presenting a system-level UWB TDOA localization framework that explicitly integrates sensor placement and evaluates its impact through real-world experiments. Using pedestrian tracking as an example, we fuse UWB TDOA and IMU measurements to estimate six-degree-of-freedom poses via an error-state Kalman filter (ESKF), and compare experimental performance against theoretical lower bounds. A challenging multi-room case study demonstrates the complete pipeline from anchor placement optimization to deployment. We additionally release our experimental dataset at http://tiny.cc/uwb_tdoa_sys_dataset to benefit the research community. To the best of our knowledge, this comprehensive evaluation, presenting a UWB TDOA-based localization system from a system-level perspective, has not been shown in the literature. A video summarizing our experiment process is available at http://tiny.cc/uwb_tdoa_sys. Our main contributions can be summarized as follows:
  • We demonstrate the accuracy and robustness of the localization system through extensive experiments and compare its performance to theoretical lower bounds.
  • We present the entire system deployment pipeline, from sensor placement optimization to system deployment and evaluation, through a challenging multi-room case study.
  • We release our experimental dataset at http://tiny.cc/uwb_tdoa_sys_dataset.
II Related Work
As UWB measurements, like any other radio frequency (RF) signals, are often affected by obstacle-induced NLOS and multi-path radio propagation, multiple approaches have been proposed to improve localization accuracy under measurement corruption. M-estimators [20] are commonly used as versatile tools to mitigate the influence of measurement outliers by employing robust cost functions to downweight the impact of large measurement residuals. Researchers have also explored novel models to represent UWB measurement residuals including both parametric [2] and non-parametric models, such as neural networks [24], Gaussian processes [10, 19], Gaussian mixture models [23], and Gaussian variational inference [17], aiming to enhance localization performance. Furthermore, continuous-time estimation techniques [7, 5] have been investigated for their applicability in asynchronous UWB-aided localization systems.
@@ II Related Work
The anchor-tag geometry, as a specific instance of the broader sensor placement problem, plays a central role in determining the performance of TDOA-based localization. In the GPS community, this effect is classically characterized using geometric dilution of precision (GDOP) [14]. Cramér-Rao lower bound (CRLB) [12] is a more general performance metric for sensor placement, which considers statistical properties of sensor measurements to evaluate estimation performance. In our previous work [22], we extended the optimal sensor placement analysis to balance the effects of anchor-tag geometry and NLOS measurement biases in cluttered indoor environments. However, our experimental validation process was limited to static experiments conducted in a laboratory setup, using multilateration as the position estimator.
In this article, we present a system-level UWB TDOA localization framework that explicitly incorporates sensor placement and evaluates its impact through real-world experiments. To achieve 6 degree-of-freedom pose estimation with a low-power, lightweight design, we fuse UWB TDOA measurements with a low-cost IMU using an error-state Kalman filter (ESKF) [13]. Through a comprehensive case study in a challenging multi-room environment, we demonstrate the full pipeline from sensor placement optimization to real-world deployment, and evaluate localization performance against corresponding theoretical analysis to provide practical design guidelines.
@@ III-B Sensor Placement Optimization
[…]
@@
Figure 4: The multi-agent pedestrian localization experiment with five agents in constellation $1$ is shown in the photo above. The estimated together with the ground truth trajectories of each agent during the experiments are shown in the bottom plots. Readers are encouraged to view our supplementary video (http://tiny.cc/uwb_tdoa_sys) to gain a better insight into our experimental process and evaluate the robustness of the localization performance.
Optimal anchor placement is essential for ensuring reliable UWB TDOA localization performance, particularly in cluttered and geometrically challenging environments. Therefore, it is important to optimize the anchor positions for regions or trajectories of interest during system design. We define a set $\mathcal{A}$ containing all possible anchor configurations with $\bm{\alpha}\in\mathcal{A}$ and use a block coordinate-wise minimization (BCM) algorithm to find the optimal placement of UWB anchors $\bm{\alpha}^{\star}$ that minimizes $\mathcal{M}(\bm{\alpha})$:
@@ III-C UWB TDOA-Inertial Navigation System
To facilitate the localization of multiple heterogeneous agents, the motion model used in this work is a 3D kinematic motion model with IMU measurements as inputs [6]. The relevant equations and a detailed description of the motion model can be found in [6]. The UWB TDOA measurements complement the motion dynamics by providing drift-free difference-of-distance measurements between UWB anchor pairs and the UWB tag. The measurement model for an anchor pair $\{\bm{a}_{i},\bm{a}_{j}\}$ is described as follows:
@@
To estimate the system state, we employ an error-state Kalman filter (ESKF) [13]. In the prediction step, inertial dead reckoning propagates the state (4) using a 3D kinematic motion model driven by IMU measurements, while the associated uncertainty is propagated accordingly. In the correction step, UWB TDOA measurements are used to estimate the error between the predicted state and the measurement-consistent state. The corresponding error state is defined as
@@
where $\delta\mathbf{p}$ and $\delta\mathbf{v}$ denote position and velocity errors, $\delta\bm{\theta}$ is the local orientation error, and $\delta\mathbf{b}_{a}$ and $\delta\mathbf{b}_{\omega}$ represent accelerometer and gyroscope bias errors, respectively. The orientation error is parameterized via a small-angle quaternion $\delta\mathbf{q}=(1,\tfrac{1}{2}\delta\bm{\theta})$, with $|\delta\bm{\theta}|\ll 1$.
The estimated error is then composed with the nominal state to correct accumulated drift, $\hat{\mathbf{x}}=\mathbf{x}\oplus\delta\mathbf{x}$, where $\oplus$ denotes quaternion multiplication for orientation and additive updates for the remaining states. Further details on the ESKF prediction and correction steps can be found in [15].
@@ III-D System Architecture and Deployment Details
To evaluate localization performance, we assembled five handheld UWB devices (Figure 3d) using the Bitcraze Loco Positioning System (LPS), which employs low-cost DWM1000 UWB radios for TDOA localization. Each device integrates a DWM1000 UWB tag, a low-cost LSM9DS1 IMU, and a Raspberry Pi 4 Model B (8 GB RAM) powered by a portable battery, all enclosed in a 3D-printed case. All system components are configured as plug-and-play to facilitate deployment. The handheld devices run Ubuntu 20.04 LTS with ROS Noetic, and the ESKF-based localization algorithm is implemented in C++, operating at $420$ Hz while consuming approximately $35\%$ of a single CPU core and $0.4\%$ of system memory.
During deployment, a millimeter-accurate Leica total station (Figure 3a) is used to establish the inertial frame and survey anchor positions. Using the “Orientate to line” procedure, the inertial frame is defined by two surveyed points, after which all anchors are measured within this frame with millimeter accuracy. The UWB localization system then provides drift-free positioning relative to the established inertial frame.
@@ IV Experiments
We evaluate the proposed UWB TDOA localization system through extensive real-world experiments across multiple environments. We first demonstrate multi-agent pedestrian localization and quantify the impact of anchor placement using two anchor constellations. We then showcase the system’s ability to provide reliable positioning under sensor occlusion. Next, we show how sensor placement analysis captures localization degradation caused by challenging anchor-tag geometry in a staircase environment. Finally, using a complex multi-room scenario consisting of a cafeteria and a narrow hallway as a case study, we demonstrate the full system pipeline—from sensor placement optimization to deployment—and compare experimental results against theoretical lower bounds. We release our experimental dataset together with development kit at http://tiny.cc/uwb_tdoa_sys_dataset. A summary video is available at http://tiny.cc/uwb_tdoa_sys.
Figure 5: Comparison of the average experimental root-mean-square error (RMSE) for multi-agent pedestrian localization in Const. 11 and 22, along with localization under occlusion in Const. 11 (Const. 1 Backpack), shown as blue bars, against the theoretical RMSE lower bounds represented by green bars.